Use the following table: $ 650,000 Big Sky Farms Partial Balance Sheet April 30, 2020 Property, plant and equipment: Land Buipplease help me with the exact calculations of this sums
Answer #1

Depreciation on building as per SLM:
Depreciation for the year ended April 30, 2021 = $975,000 / 15 = $65,000

Depreciation for Equipment under Double Declining balance method:
Depreciation as per SLM = ($750,000 - $250,000) / 10 years = $50,000
Depreciation rate SLM = $50,000 / ($750,000 - $250,000) = 10%
Dep. rate as per DDB = 10% * 2 = 20%

Depreciation Schedule Year Book Value Depreciation Depreciation Accumulated Book Value Year Start Percent Expense Depreciatio

Date Account Titles Debit Credit
April 30, 2021 Depreciation expense, Building $65,000
   Accumulated Depreciation $65,000
April 30, 2021 Depreciation expense, Equipment $345,600
Accumulated Depreciation $345,600